How many grams of H 2 O could be produced from the reaction of 70.8 grams of Mg(OH) 2 ? Given the equation , Mg(OH) 2 (aq)+2 HCl(aq) MgCl 2 (aq)+2 H 2 O(l)
1
Expert's answer
2020-04-22T09:18:48-0400

mole ratio=Mg(OH)2: H2O= 1:2

"moles= mass\/RMM"

If 70.8g of Mg(OH)2 react then 70.8/58.3=1.214 moles reacted

therefore 2.428 moles of water was produced.

mass of water=2.428*18=43.7g